OpenStudy (anonymous):

3sqrt10+7 sqrt 15-6 sqrt 10-4 sqrt 15

OpenStudy (jiteshmeghwal9):

combine like terms\[3\sqrt{10}-6\sqrt{10}+7\sqrt{15}-4\sqrt{15}\]\[-3\sqrt{10}+3\sqrt{15}\]